OpenStudy (elenathehomeschooler):

Help! 12. A bacteria culture has 2,000 bacteria. The number of bacteria increases by 5% each hour. How many bacteria are there after 12 hours?

OpenStudy (elenathehomeschooler):

@jim_thompson5910 can you help??

jimthompson5910 (jim_thompson5910):

From the word problem, the equation that forms is y = 2000*(1.05)^x where x = number of hours y = amount of bacteria

jimthompson5910 (jim_thompson5910):

plug in x = 12 and tell me what you get for y

OpenStudy (elenathehomeschooler):

y= 3591.7 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

2000*(1.05)^12 = 3,591.71265204427 it looks good so roughly 3,591 bacteria

OpenStudy (elenathehomeschooler):

Okay thanks!